# STARFormer: A Novel Spatio-Temporal Aggregation Reorganization Transformer of FMRI for Brain Disorder Diagnosis

### v1

@@ Abstract
Many existing methods that use functional magnetic resonance imaging (fMRI) classify brain disorders, such as autism spectrum disorder (ASD) and attention deficit hyperactivity disorder (ADHD), often overlook the integration of spatial and temporal dependencies of the blood oxygen level-dependent (BOLD) signals, which may lead to inaccurate or imprecise classification results. To solve this problem, we propose a Spatio-Temporal Aggregation eorganization ransformer (STARFormer) that effectively captures both spatial and temporal features of BOLD signals by incorporating three key modules. The region of interest (ROI) spatial structure analysis module uses eigenvector centrality (EC) to reorganize brain regions based on effective connectivity, highlighting critical spatial relationships relevant to the brain disorder. The temporal feature reorganization module systematically segments the time series into equal-dimensional window tokens and captures multiscale features through variable window and cross-window attention. The spatio-temporal feature fusion module employs a parallel transformer architecture with dedicated temporal and spatial branches to extract integrated features. The proposed STARFormer has been rigorously evaluated on two publicly available datasets for the classification of ASD and ADHD. The experimental results confirm that the STARFormer achieves state-of-the-art performance across multiple evaluation metrics, providing a more accurate and reliable tool for the diagnosis of brain disorders and biomedical research. The codes will be available at: https://github.com/NZWANG/STARFormer.
{IEEEkeywords}
Brain disorder diagnosis, fMRI, eigenvector centrality, spatio-temporal information integration, transformer.
@@ 1 Introduction
The human brain function can be characterized by intricate functional networks where multiple brain regions cooperate to facilitate cognitive processes and mental states [1]. Disruptions of these networks can manifest as various neurodevelopmental conditions, such as autism spectrum disorder (ASD) [2] and attention deficit hyperactivity disorder (ADHD) [3]. However, the underlying mechanisms of these disruptions are not yet fully understood [4, 5, 6]. Resting-state functional magnetic resonance imaging (rs-fMRI) has emerged as a powerful, noninvasive technique for investigating these functional networks by measuring blood oxygen-level-dependent (BOLD) signals with relatively high spatial and temporal resolution [7]. In particular, functional connectivity (FC), which reflects the temporal correlations of BOLD signals between regions of interest (ROI), provides crucial insights into brain organization and cognitive functions [8, 9].
@@
Traditional diagnostic approaches, which primarily rely on symptomatic observations and clinical expertise, are inadequate for detecting intricate patterns across the entire brain. This limitation highlights the need for more advanced analytical methods. Initially, traditional machine learning techniques were used to analyze multivariate brain responses from rs-fMRI data for diagnosing neurodevelopmental disorders [10]. The field then progressed with the advent of deep learning [11] and, more recently, the groundbreaking introduction of transformer models [12]. Transformers have significantly improved the modeling of complex patterns in high-dimensional data through self-attention mechanisms, offering better scalability and more efficient capture of global network information than conventional approaches [13, 14]. However, traditional transformer models often struggle to capture both the intricate spatial and temporal features within fMRI data simultaneously. The standard self-attention mechanism in transformer encoders tends to focus on identifying parts of the time series with similar patterns across the entire sequence, especially those with matching peaks. This approach may overlook patterns that are similar in the short to medium term, as they occur closely in time. In fMRI data analysis, recognizing these short- to medium-term patterns is crucial because significant changes in BOLD signals may only occur within shorter time windows, rather than being consistent throughout the entire duration [15].
@@
To address these limitations, we propose a novel spatio-temporal aggregation reorganization transformer (STARFormer) that effectively captures both spatial and temporal information of brain functional networks through a designed dual-branch architecture. This approach uses effective connectivity to provide spatial information for the time series of BOLD signals. Unlike methods that rely solely on FC, this approach preserves temporal dynamics while incorporating spatio-temporal information. Our methodology introduces three key innovations:
@@
1) ROI Spatial Structure Analysis Module: This module employs eigenvector centrality (EC) to reorganize brain regions based on their functional importance within seven established brain networks, ensuring the preservation and enhancement of crucial spatial relationships. This reorganization significantly improves the model’s ability to identify disorder-related spatial patterns.
@@ 2.1 Centrality Nodes Identification in Brain Disorders
Identifying critical brain nodes is crucial for understanding disorder-specific functional disruptions, as neurological disorders significantly differ in regional connectivity patterns [16]. For example, individuals with mild cognitive impairment exhibit altered effective connectivity in memory-related areas such as the hippocampus and amygdala [17], while individuals with ASD display distinctive alterations in regions associated with social cognition, such as the default mode network [18].
Measures of node centrality have been effectively applied in diagnosing brain disorders to uncover differential patterns. For example, Saha [19] used EC to assess group differences in the centrality of brain regions, distinguishing between typically developing children and those with ASD. Similarly, Grobelny et al. [20] utilized betweenness centrality (BC) to construct diagnostic models for pediatric epilepsy, revealing that nodes with high BC during seizure onset could represent centers in self-regulating networks that help terminate seizures. Liao et al. [21] found that changes in degree centrality (DC) in Parkinson’s disease were frequency-related and frequency-specific. However, due to computational complexity, BC, DC and other centrality measurement methods are less suitable for analyzing full-brain networks that involve a large number of voxels [19]. Hence, this study focuses on EC to identify differential connectivity patterns, highlighting the importance of key nodes in distinguishing pathological alterations.
@@ 2.2 Features of fMRI for Brain Disorders Identification
The extraction of FC features is a key focus in many neuroimaging studies. A typical approach for identifying brain disorders involves extracting FC, which represents the temporal correlation matrix of BOLD signals from ROI, followed by the use of classifiers such as SVM and logistic regression to reduce dimensionality [10]. A study by [22] proposed a novel architecture called local sequential feature coupling global representation learning (LCGNet), which uses convolution operations and self-attention mechanisms to enhance representation learning in fully convolutional networks for automatic brain disorder classification. However, FC is limited to capturing linear relationships and lacks causal or directional insights. The effective connectivity approach has emerged as a promising tool that describes brain activity by incorporating causal interactions between brain regions [23]. The study by Dai et al. [24] indicated that patients with major depressive disorder (MMD) exhibit significantly altered effective connectivity networks in various brain networks during the resting state, which may serve as potential biomarkers. Unlike FC, which focuses solely on statistical correlations, effective connectivity elucidates the directional influence between regions, providing a deeper understanding of the connections within the brain [25].
Although deep learning methods based on FC have been widely used for diagnosing brain diorders, some studies reveal that these methods can overlook the temporal nature of fMRI data, thereby losing vital information on temporal dynamics [26]. Recent advances in time-domain-based deep learning methods have shown great promise in diagnosing brain disorders. For instance, long short-term memory (LSTM) networks have been used to classify individuals with ASD and typical controls from multi-site fMRI time series [27], while a transformer model with a fusion window has further advanced fMRI time series analysis [28].
@@ 2.3 Spatio-Temporal Information Integration for Brain Disorders Identification
For modeling fMRI features, most deep learning models do not fully exploit the spatial and temporal dependencies of fMRI signals, limiting the precise analysis of brain activity. The gradient-weighted Markov random field (gwMRF) model mentioned in [29] enables spatially-ordered brain region extraction. This study demonstrated that spatial disorganization decreases model performance, emphasizing the need to preserve spatial information in rs-fMRI. Recent studies increasingly emphasize the importance of spatial-temporal information integration. For example, the use of 3D-CNNs can extract spatial features from fMRI data for diagnosing ASD and ADHD [30, 31]. Liu et al. [32] proposed STCAL, a spatio-temporal cooperative attention learning model that employs a guided cooperative attention module to simultaneously capture spatio-temporal correlations and learn fine-grained attention representations from time series fMRI data. Zhang et al. [33] used independent component analysis (ICA) to aggregate spatio-temporal information from fMRI, improving the performance of depression diagnosis.
Transformer models have also made significant advances in integrating spatio-temporal features, showcasing their potential to capture complex dependencies within fMRI data. In [26], a novel transformer-based framework, the ST-Transformer, was introduced, featuring a linear spatio-temporal multihead attention unit to extract spatio-temporal features from fMRI data through spatial self-attention. Similarly, a spatio-temporal graph transformer network was proposed in [34], incorporating a spatial transformer-based graph message-passing mechanism to capture inter-regional relationships and using FC as edge features.
@@ 3.1 ROI Spatial Structure Analysis Module
The spatial information of the brain is provided by the effective connectivity matrix $\mathbf{G}$ of N patients. We use Granger causality (GC) to calculate effective connectivity between different brain regions, as it can reveal whether the time series of a brain region predict the time series of another [35]. Suppose the chosen atlas divides the brain into $n$ ROIs, each with a time series $s_{i}(t)$ of length $m$. We first establish an autoregressive model, where the time series of each ROI $i$ can be predicted using its own values from the past $h$ time points:
@@
The ROIs and the connectivity metrics of the effective connectivity matrix $\mathbf{G}$ define the nodes and edges of the graph, respectively. Serving as the adjacency matrix of the network, $\mathbf{G}$ facilitates the assessment of node centrality through its eigenvectors. EC quantifies the importance of the nodes within the network by considering not only the number of connections each node possesses, but also the importance of the nodes to which it is connected. The EC score $p_{i}$ for each ROI $i$ measures its connection strength to other influential ROIs in the network and can be expressed as
@@
where $\mathbf{P}=[p_{1},p_{2},...,p_{n}]^{T}$ represents the EC vector, with $\lambda$ denoting its corresponding eigenvalue. The calculation requires determining the principal eigenvector $\mathbf{P}$ of the adjacency matrix $\mathbf{G}$ in conjunction with its dominant eigenvalue $\lambda$. Subsequently, the EC of any given ROI $i$ is precisely captured by the component $p_{i}$ within this principal eigenvector associated with the maximal eigenvalue $\lambda$. To guarantee uniqueness and stability, $p_{i}$ is normalized such that
@@
Following the computation of the EC vector for ROIs of each patient, the average EC vector $\mathbf{\bar{P}}$ for all patients was determined by
@@
We consider using the obtained EC score $\bar{p_{i}}$ of ROI to rearrange the spatial information. To reduce noise and errors caused by unrelated regional arrangements, ROIs are grouped according to the seven functional networks of the brain (visual network, somatomotor network, default network, limbic network, dorsal attention network, ventral attention and frontoparietal network) [36]. This functional network-based partition reflects the modular and distributed nature of brain function, with different functional networks responsible for processing different types of information and tasks. The ROIs within each group are then arranged by $\bar{p_{i}}$ in descending order:
@@
The global EC of ROIs allows for better capture of interactions or dependencies between regions. Rearranging ROIs aligns the order of the brain network with the spatial characteristics of the patients, improving the stability and consistency of its analysis. The sorting results of the ROIs are applied to the fMRI time series data for all participants $S=\{S_{1}(t),S_{2}(t),...,S_{n}(t)\}$, and
@@ 3.2 Spatio-Temporal Feature Fusion Module
Arrayed rs-fMRI data $S^{\prime}$ are used as input to the spatio-temporal feature fusion module. $S^{\prime}$ enter the temporal branch and the spatial branch, respectively. The temporal branch treats each row $S^{\prime}_{i}(t)$ in $S^{\prime}$ as a token, allowing the cross-window attention in the temporal branch to learn local features. When inputting into the spatial branch, the spatial branch adjusts the dimensions of $S^{\prime}$ to $m\times n$ as the new input $S^{T}=\{t_{1},t_{2},...,t_{m}\}$, treating each column $t_{i}$ in $S^{\prime}$ as a token. Self-attention in the spatial branch can learn the global dependencies between brain regions. The results obtained from the two branches are then integrated, and the detailed operations can be expressed as
@@
where $Sp$ refers to the processing of the spatial branch and $Te$ refers to the processing of the temporal branch. Finally, the combined feature $F$ was input into the multilayer perceptron (MLP) for classification.

### v2

@@ Abstract
Many existing methods that use functional magnetic resonance imaging (fMRI) to classify brain disorders, such as autism spectrum disorder (ASD) and attention deficit hyperactivity disorder (ADHD), often overlook the integration of spatial and temporal dependencies of the blood oxygen level-dependent (BOLD) signals, which may lead to inaccurate or imprecise classification results. To solve this problem, we propose a Spatio-Temporal Aggregation Reorganization Transformer (STARFormer) that effectively captures both spatial and temporal features of BOLD signals by incorporating three key modules. The region of interest (ROI) spatial structure analysis module uses eigenvector centrality (EC) to reorganize brain regions based on effective connectivity, highlighting critical spatial relationships relevant to the brain disorder. The temporal feature reorganization module systematically segments the time series into equal-dimensional window tokens and captures multiscale features through variable window and cross-window attention. The spatio-temporal feature fusion module employs a parallel transformer architecture with dedicated temporal and spatial branches to extract integrated features. The proposed STARFormer has been rigorously evaluated on two publicly available datasets for the classification of ASD and ADHD. The experimental results confirm that STARFormer achieves state-of-the-art performance across multiple evaluation metrics, providing a more accurate and reliable tool for the diagnosis of brain disorders and biomedical research. The official implementation codes are available at: https://github.com/NZWANG/STARFormer.
Keywords Brain disorder diagnosis, fMRI, eigenvector centrality, spatio-temporal information integration, transformer
@@ 1 Introduction
The human brain function can be characterized by intricate functional networks where multiple brain regions cooperate to facilitate cognitive processes and mental states (zhang2024learning, ). Disruptions of these networks can manifest as various neurodevelopmental conditions, such as autism spectrum disorder (ASD) (bicks2024functional, ) and attention deficit hyperactivity disorder (ADHD) (koirala2024neurobiology, ). However, the underlying mechanisms of these disruptions are not yet fully understood (pievani2011functional, ; wee2014disrupted, ; qian2019large, ). Resting-state functional magnetic resonance imaging (rs-fMRI) has emerged as a powerful, noninvasive technique for investigating these functional networks by measuring blood oxygen-level-dependent (BOLD) signals with relatively high spatial and temporal resolution (Pei2025bold, ). In particular, functional connectivity (FC), which reflects the temporal correlations of BOLD signals between regions of interest (ROI), provides crucial insights into brain organization and cognitive functions (li2024mhnet, ; Zhou2024fusing, ).
@@
Traditional diagnostic approaches, which primarily rely on symptomatic observations and clinical expertise, are inadequate for detecting intricate patterns across the entire brain. This limitation highlights the need for more advanced analytical methods. Initially, traditional machine learning techniques were used to analyze multivariate brain responses from rs-fMRI data for diagnosing neurodevelopmental disorders (khosla2019machine, ). The field then progressed with the advent of deep learning (shoeibi2023diagnosis, ) and, more recently, the groundbreaking introduction of transformer models (vaswani2017attention, ). Transformers have significantly improved the modeling of complex patterns in high-dimensional data through self-attention mechanisms, offering better scalability and more efficient capture of global network information than conventional approaches (cong2024comprehensive, ; Mubonanyikuzo2024DetectionOA, ). However, traditional transformer models often struggle to capture both the intricate spatial and temporal features within fMRI data simultaneously. The standard self-attention mechanism in transformer encoders tends to focus on identifying parts of the time series with similar patterns across the entire sequence, especially those with matching peaks. This approach may overlook patterns that are similar in the short to medium term, as they occur closely in time. In fMRI data analysis, recognizing these short- to medium-term patterns is crucial because significant changes in BOLD signals may only occur within shorter time windows, rather than being consistent throughout the entire duration (hutchison2013dynamic, ).
@@
To address these limitations, we propose a novel spatio-temporal aggregation reorganization transformer (STARFormer) that effectively captures both spatial and temporal information of brain functional networks through a designed dual-branch architecture. This approach uses effective connectivity to provide spatial information for the time series of BOLD signals. Unlike methods that rely solely on FC, this approach preserves temporal dynamics while incorporating spatio-temporal information. STARFormer, through ROI rearrangement and a variable window strategy, not only improves diagnostic accuracy but also reduces computational complexity by leveraging windowed computations. Compared with convolution-attention hybrid architecture of LCGNet (zhou2024lcgnet, ), dual-branch design of STARFormer separates and integrates spatiotemporal features more effectively, while providing stronger interpretability via ROI analysis. These improvements allow STARFormer to achieve more accurate diagnosis and better clinical interpretability while maintaining computational efficiency.
For spatial information modeling, we specifically choose eigenvector centrality (EC) over other centrality measures based on several critical considerations (EC, ). First, EC measures the global influence of a node within the entire network-it depends not only on the node’s direct connections but also on the centrality of the nodes it connects to. Second, the Transformer architecture inherently lacks spatial structure inductive bias and requires externally provided ordering or positional information to learn the organizational patterns between nodes. EC not only identifies globally important nodes in brain networks but also provides a stable and coherent spatial ordering method. This biologically grounded approach effectively incorporates the network roles of ROIs into the input structure, guiding the Transformer to focus on key brain regions and thereby enhancing brain disorder classification performance.
Our methodology introduces three key innovations:
@@
1) ROI Spatial Structure Analysis Module: This module employs EC to reorganize brain regions based on their functional importance within seven established brain networks, ensuring the preservation and enhancement of crucial spatial relationships. This reorganization significantly improves the model’s ability to identify disorder-related spatial patterns.
@@ 2.1 Centrality Nodes Identification in Brain Disorders
Identifying critical brain nodes is crucial for understanding disorder-specific functional disruptions, as neurological disorders significantly differ in regional connectivity patterns (van2019cross, ). For example, individuals with mild cognitive impairment exhibit altered effective connectivity in memory-related areas such as the hippocampus and amygdala (zheng2017altered, ), while individuals with ASD display distinctive alterations in regions associated with social cognition, such as the default mode network (monk2009abnormalities, ).
Measures of node centrality have been effectively applied in diagnosing brain disorders to uncover differential patterns. For example, Saha (saha2024eigenvector, ) used EC to assess group differences in the centrality of brain regions, distinguishing between typically developing children and those with ASD. Similarly, Grobelny et al. (grobelny2018betweenness, ) utilized betweenness centrality (BC) to construct diagnostic models for pediatric epilepsy, revealing that nodes with high BC during seizure onset could represent centers in self-regulating networks that help terminate seizures. Liao et al. (liao2021changes, ) found that changes in degree centrality (DC) in Parkinson’s disease were frequency-related and frequency-specific.
However, due to computational complexity, BC, DC and other centrality measurement methods are less suitable for analyzing full-brain networks that involve a large number of voxels (saha2024eigenvector, ). Hence, this study focuses on EC to identify differential connectivity patterns, highlighting the importance of key nodes in distinguishing pathological alterations.
@@ 2.2 Features of fMRI for Brain Disorders Identification
Besides directly exploring the relationship between node centrality and brain disorders, the extraction of FC features is also a key focus in many brain disorder identification studies. A typical approach for identifying brain disorders involves extracting FC, which represents the temporal correlation matrix of BOLD signals from ROI, followed by the use of classifiers such as SVM and logistic regression to reduce dimensionality (khosla2019machine, ). A study by (zhou2024lcgnet, ) proposed a novel architecture called local sequential feature coupling global representation learning (LCGNet), which uses convolution operations and self-attention mechanisms to enhance representation learning in fully convolutional networks for automatic brain disorder classification.
However, FC is limited to capturing linear relationships and lacks causal or directional insights. The effective connectivity approach has emerged as a promising tool that describes brain activity by incorporating causal interactions between brain regions (wang2024deep, ). The study by Dai et al. (dai2023altered, ) indicated that patients with major depressive disorder (MMD) exhibit significantly altered effective connectivity networks in various brain networks during the resting state, which may serve as potential biomarkers. Unlike FC, which focuses solely on statistical correlations, effective connectivity elucidates the directional influence between regions, providing a deeper understanding of the connections within the brain (sharaev2016effective, ).
@@
Although deep learning methods based on FC have been widely used for diagnosing brain diorders, some studies reveal that these methods can overlook the temporal nature of fMRI data, thereby losing vital information on temporal dynamics (deng2022sttransformer, ). Recent advances in time-domain-based deep learning methods have shown great promise in diagnosing brain disorders. For instance, long short-term memory (LSTM) networks have been used to classify individuals with ASD and typical controls from multi-site fMRI time series (dvornek2017lstm, ), while a transformer model with a fusion window has further advanced fMRI time series analysis (bedel2023bolt, ).
@@ 2.3 Spatio-Temporal Information Integration for Brain Disorders Identification
For modeling fMRI features, most deep learning models use features like FC but do not fully exploit the spatial and temporal dependencies of fMRI signals, which limits the precise analysis of brain activity. The gradient-weighted Markov random field (gwMRF) model mentioned in (wang2024multi, ) enables spatially-ordered brain region extraction. This study demonstrated that spatial disorganization decreases model performance, emphasizing the need to preserve spatial information in rs-fMRI. Recent studies increasingly emphasize the importance of spatial-temporal information integration. For example, the use of 3D-CNNs can extract spatial features from fMRI data for diagnosing ASD and ADHD (li20182, ; zou20173d, ). Liu et al. (liu2023stcal, ) proposed STCAL, a spatio-temporal cooperative attention learning model that employs a guided cooperative attention module to simultaneously capture spatio-temporal correlations and learn fine-grained attention representations from time series fMRI data. Zhang et al. (tomography10120138, ) used independent component analysis (ICA) to aggregate spatio-temporal information from fMRI, improving the performance of depression diagnosis.
Transformer models have also made significant advances in integrating spatio-temporal features, showcasing their potential to capture complex dependencies within fMRI data. In (deng2022sttransformer, ), a novel transformer-based framework, the ST-Transformer, was introduced, featuring a linear spatio-temporal multihead attention unit to extract spatio-temporal features from fMRI data through spatial self-attention. Similarly, a spatio-temporal graph transformer network was proposed in (he2024stgtn, ), incorporating a spatial transformer-based graph message-passing mechanism to capture inter-regional relationships and using FC as edge features.
Although existing studies have advanced brain disorder diagnosis, they exhibit critical limitations. Many methods ignore temporal dynamics of fMRI signals or overlook spatial structure relationships, while few provide biologically interpretable explanations for clinical adoption. To address these limitations, STARFormer directly utilizes fMRI time-series data to capture temporal dynamics, integrates spatial relationships through EC-based spatial feature reorganization to enhance both spatial structure learning and biological interpretability for more accurate and clinically relevant diagnosis.
@@ 3.1 ROI Spatial Structure Analysis Module
The spatial information of the brain is provided by the effective connectivity matrix $\mathbf{G}$ of N patients. We use Granger causality (GC) to calculate effective connectivity between different brain regions, as it can reveal whether the time series of a brain region predict the time series of another (shojaie2022granger, ). Suppose the chosen atlas divides the brain into $n$ ROIs, each with a time series $s_{i}(t)$ of length $m$. We first establish an autoregressive model, where the time series of each ROI $i$ can be predicted using its own values from the past $h$ time points:
@@
The ROIs and the connectivity metrics of the effective connectivity matrix $\mathbf{G}$ define the nodes and edges of the graph, respectively. Serving as the adjacency matrix of the network, $\mathbf{G}$ facilitates the assessment of node centrality through its eigenvectors.
We chose EC to identify key brain regions because it more accurately captures global influence while remaining computationally efficient and theoretically aligned. Unlike DC, which merely counts direct connections, EC recursively weights each link by the importance of its neighbors to reflect a node’s true influence across the entire network. In contrast to BC-whose O(n³) complexity is impractical for full-brain networks of hundreds of ROIs—and closeness centrality(CC)-which has theoretical limitations in directed networks—EC remains both efficient and robust. Crucially, its iterative calculation naturally models the propagation of causal influences, making it ideal for pinpointing hub regions whose altered connectivity underlies disease.
The EC score $p_{i}$ for each ROI $i$ measures its connection strength to other influential ROIs in the network and can be expressed as
@@
where $\mathbf{P}=[p_{1},p_{2},...,p_{n}]^{T}\in R^{n\times 1}$ represents the EC vector, with $\lambda$ denoting its corresponding eigenvalue. The calculation requires determining the principal eigenvector $\mathbf{P}$ of the adjacency matrix $\mathbf{G}$ in conjunction with its dominant eigenvalue $\lambda$. Subsequently, the EC of any given ROI $i$ is precisely captured by the component $p_{i}$ within this principal eigenvector associated with the maximal eigenvalue $\lambda$. To guarantee uniqueness and stability, $p_{i}$ is normalized such that
@@
Following the computation of the EC vector for ROIs of each patient, the average EC vector $\mathbf{\bar{P}}\in R^{n\times 1}$ for all patients was determined by
@@
We consider using the obtained EC score $\bar{p_{i}}$ of ROI to rearrange the spatial information. To reduce noise and errors caused by unrelated regional arrangements, ROIs are grouped according to the seven functional networks of the brain (visual network, somatomotor network, default network, limbic network, dorsal attention network, ventral attention and frontoparietal network) (yeo2011organization, ). A hierarchical ranking strategy not only prevents functionally related regions from being scattered but also captures disease-specific important ROIs, producing an ROI sequence that is more robust to disease variability and noise. The ROIs within each group are then arranged by $\bar{p_{i}}$ in descending order:
@@
The global EC of ROIs allows for better capture of interactions or dependencies between regions. Rearranging ROIs aligns the order of the brain network with the spatial characteristics of the patients, improving the stability and consistency of its analysis. The sorting results of the ROIs are applied to the fMRI time series data for all participants $S=\{S_{1}(t),S_{2}(t),...,S_{n}(t)\}\in R^{n\times m}$, and
@@ 3.2 Spatio-Temporal Feature Fusion Module
Arrayed rs-fMRI data $S^{\prime}$ are used as input to the spatio-temporal feature fusion module. $S^{\prime}$ enter the temporal branch and the spatial branch, respectively. The temporal branch treats each row $S^{\prime}_{i}(t)$ in $S^{\prime}$ as a token, allowing the cross-window attention in the temporal branch to learn local features. When inputting into the spatial branch, the spatial branch adjusts the dimensions of $S^{\prime}$ to $m\times n$ as the new input $S^{T}=\{t_{1},t_{2},...,t_{m}\}\in R^{m\times n}$, treating each column $t_{i}$ in $S^{\prime}$ as a token. Self-attention in the spatial branch can learn the global dependencies between brain regions. The results obtained from the two branches are then integrated, and the detailed operations can be expressed as
@@
where $Sp$ refers to the processing of the spatial branch and $Te$ refers to the processing of the temporal branch. Finally, the combined feature $F$ was input into the multilayer perceptron (MLP) for classification. The MLP has two layers, uses ReLU activation, and a hidden dimension of 256.